# Unifying Appearance Codes and Bilateral Grids for Driving Scene Gaussian Splatting

### v1

@@ 4.1 Experimental Setup
Datasets.  We evaluate on four autonomous driving datasets: Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset , selected for their diversity in sensor configurations (LiDAR, camera specifications), environmental conditions (lighting, seasons), and geographic locations. Identical model hyperparameters are used across all datasets to test generalization capability. Technical specifications including camera counts and ego-vehicle view cropping details are provided in Appendix A2.
Baseline.  Our focus is on modeling appearance variance across viewpoints, we build upon the open-source works DriveStudio chen2024omnire , ChatSim wei2024editable , StreetGS yan2024street and test three approaches: appearance codes, standalone bilateral grids, and our multi-scale bilateral grids. This allows for direct comparison of how each method handles photometric inconsistencies across viewpoints.
@@
Table 2: Comparison with different baseline approaches for scene reconstruction. Performance is reported as the average over the nuScenes caesar2020nuscenes scenarios.
Implementation details.  We evaluate our method on many autonomous driving datasets to demonstrate its robustness. For each dataset, We train all the methods on a single NVIDIA L40 GPU and compare the 3D reconstruction and novel view synthesis results between camera and LiDAR data. Additionally, we ablate key components of our method and quantify their impact on both appearance and geometry quality.
Dataset Method PSNR $\uparrow$
Full Image Human Vehicle
Waymo sun2020scalability Baseline 26.40 23.91 21.25
Baseline w/AC 26.44 23.81 21.34
Baseline w/BG 21.51 22.91 19.82
Ours 26.55 23.97 21.36
NuScenes caesar2020nuscenes Baseline 23.74 23.55 22.80
Baseline w/AC 23.74 23.35 22.84
Baseline w/BG 23.60 23.63 22.79
Ours 24.64 23.75 23.22
Argoverse wilson2023argoverse Baseline 22.53 20.73 20.84
Baseline w/AC 22.51 20.70 20.80
Baseline w/BG 21.70 20.67 20.59
Ours 22.58 20.79 20.86
Pandaset xiao2021pandaset Baseline 27.49 26.94 21.02
Baseline w/AC 27.51 26.98 21.13
Baseline w/BG 27.38 27.08 20.77
Ours 27.89 27.22 21.32
Table 3: Comparison on Novel View Synthesis.
@@
Furthermore, we introduce two additional regularization terms to enhance image fidelity: Adaptive Total Variation Regularization. This term encourages smoothness and reduces noise while preserving image details, and Circle Regularization Loss. This loss applies an inverse transformation to the ground-truth images, preventing discrepancies and image quality degradation. As visually demonstrated in Fig. 5, these two terms significantly improve image fidelity. Detailed explanations are provided in Appendix A1.
2) Rendering. For dynamic rendering using our multi-scale bilateral grid framework, we employ a specific interpolation strategy when encountering novel test images, especially those simulating dynamic ISP conditions. Our proposed method first conducts a temporal proximity search to identify the most relevant grids. Following this, we perform scale-specific interpolation using the two nearest grids found. More comprehensive details regarding this interpolation approach are available in Appendix A1.
Figure 3: Visualizations of our method and baseline methods on Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset dataset.
Figure 4: Our proposed framework (c) outperforms appearance codes (a) and single bilateral grids (b) by simultaneously addressing optimization challenges and enhancing geometric modeling. This delivers significant geometric accuracy improvements (lower Chamfer Distance (CD) via multi-scale bilateral grid, reducing error and improving reconstruction) and reduced photometric inconsistency (fewer floaters) in dynamic, decoupled driving scenes. Yellow indicates a high lidar error, while Purple indicates a low lidar error.
@@ 4.2 Quantitative Evaluation
Reconstruction Quality.  We evaluate the reconstruction quality of our method using appearance and geometry metrics, with detailed results presented in Tab. 1.
(1) Appearance Fidelity. Our approach demonstrates superior performance in appearance-based metrics (PSNR and SSIM). As shown in Tab. 1, our method achieves the highest PSNR and SSIM scores for full image reconstruction across all tested datasets (Waymo, NuScenes, Argoverse, and Pandaset). Furthermore, when evaluating segmented dynamic regions such as humans and vehicles, our method largely surpasses the baselines, indicating its robustness in capturing both global scene appearance and fine-grained details of dynamic objects. For instance, on NuScenes, our PSNR for ’Vehicle’ is 27.31, compared to the best baseline score of 26.52. While there are isolated cases where a baseline performs comparably or slightly better (e.g., Human SSIM on Argoverse), our method generally offers the most consistent and highest quality appearance reconstruction.
(2) Geometric Accuracy. In terms of geometric fidelity, our method consistently outperforms all baselines across all datasets, as evidenced by the Chamfer Distance (CD), Root Mean Square Error (RMSE), and Depth L2 Error metrics in Tab. 1. For example, on the Waymo dataset, our method achieves a CD of 0.989, significantly lower than the best baseline score of 1.378. This superior geometric accuracy is partly attributed to our model’s improved handling of "floater" artifacts, leading to more precise LiDAR representations, as qualitatively illustrated in Fig. 4.
Novel View Synthesis.  The performance of our method in novel view synthesis is detailed in Tab. 3. Our approach consistently achieves the highest PSNR scores across all datasets for full images as well as for segmented human and vehicle categories. For example, on the NuScenes dataset, our method achieves a PSNR of 24.64 for the full image, 23.75 for humans, and 23.22 for vehicles, outperforming all baseline variations. This underscores the capability of our method to generate high-fidelity novel views from new camera poses.
Comparison with State-of-the-Art Methods. To further assess the benefits of our architectural contributions, we integrated our core components into two recent state-of-the-art methods: ChatSim wei2024editable and StreetGS yan2024street . The comparative results, averaged over NuScenes scenarios, are presented in Tab. 2.
When our enhancements are applied (denoted as "ChatSim (Ours)" and "StreetGS (Ours)"), both methods exhibit significant improvements across all evaluated aspects. For instance, integrating our approach with ChatSim improves its reconstruction PSNR from 25.10 to 27.04 and its novel view synthesis PSNR from 23.27 to 24.67. Similarly, StreetGS sees its reconstruction PSNR boosted from 25.74 to 27.90 and its Chamfer Distance reduced from 1.604 to 1.272. These results highlight the generalizability and effectiveness of our proposed techniques, demonstrating their capability to enhance the performance of existing strong baseline models for comprehensive scene understanding and representation.
@@ 4.3 Qualitative Evaluation
Rendering Results.  Fig. 3 presents visualization results of our method on scenes from Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset , demonstrating its capability to handle diverse conditions that might otherwise lead to inconsistencies.
Figure 5: Visualizations of ablations on Circle Regularization loss and Adaptive TV loss.
Figure 6: Visualization of Affine Transformations and Distributions. The first three rows display the affine transformation results and corresponding distributions when applying the first level (scale 0), the first two levels (scales 0-1), and all three levels (scales 0-1-2) of our multi-scale approach. The final row compares these to results from a single bilateral grid, demonstrating that the multi-scale bilateral grid achieves a coarse-to-fine, smooth image enhancement with more diverse processing capabilities.
Affine Transformations Distributions.  According to Fig. 6, analyzing grayscale histograms of learned affine transformations reveals a key distinction between original single-scale and our multi-scale bilateral grids. For the original grid, each viewpoint’s histogram typically exhibits two distinct peaks. This bi-modal distribution suggests a focused, somewhat limited photometric adaptation. It implies that for each view, correction focuses on a few dominant transformations. More visualization results are shown at Fig. 3. A particularly interesting observation arises examining histograms of our multi-scale grid’s coarsest level. Intriguingly, these histograms also demonstrate peaky distributions, varying significantly across viewpoints. This isn’t contradictory, but insightful: even at its coarsest level, the grid learns view-dependent base appearances. The coarse grid captures dominant, scene-wide photometric transformations characteristic of each viewpoint. Peaks in coarse-level histograms likely represent these primary, global photometric adjustments, tailored per view.
@@
Table 4: Computational Efficiency Analysis. †  means using bilateral grids processing during rendering.
However, the key advantage is evident in the overlaid histogram, which aggregates distributions from all multi-scale levels and viewpoints. This histogram is notably flatter and more dispersed than those of the original grid, a characteristic directly resulting from hierarchical and residual refinement. Our framework first uses the coarse grid to establish a view-dependent base appearance. Subsequently, medium and fine grids learn residual transformations to correct coarse grid errors. Medium grids then refine regional photometric variations, while fine grids handle local, pixel-level details. This layered, residual approach significantly broadens the spectrum of learnable photometric transformations. The original grid’s two-peaked histograms signify its narrow adaptation. In contrast, the multi-scale version’s flatter histogram reveals its core strength: a diverse, generalizable, and robust representation of spatially-varying photometric transformations. This enhanced representation is crucial for consistent and high-quality renderings across diverse driving views.
@@ 4.4 Ablation Studies
Table 5: Ablations. Numbers in Guidance Map Resolution settings indicate downsample factors from coarse level to fine level, and numbers in Grid Size Combinations settings indicate grid size.
Guidance Map Resolution.   Varying guidance map resolution shows a trade-off in performance. Using the highest resolution (1,1,1) results in a slightly lower NVS PSNR compared to using (2,2,1) which achieves the highest NVS PSNR(24.64) in these ablations, demonstrating our patch-based approach balances efficiency and effectiveness. Grid Size Combinations.   Using different combinations of grid sizes significantly impacts geometry, as indicated by CD. Combining smaller grids like (2,2,1)+(4,4,2) results in a lower CD (1.213) than using combinations including larger grids or the combination with the largest grid (4,4,2)+(8,8,4)+(16,16,8). Loss Functions.   Ablating the circle regularization loss slightly reduces both Reconstruction PSNR (from 27.69 to 27.47) and NVS PSNR (from 24.64 to 24.52). Removing the adaptive total variance loss has a minimal impact on rendering result. Efficiency.   As shown in Tab. 4, while the ’Baseline w/BG’ dramatically increases parameters and training time, our method achieves a much more moderate increase in parameters and training time (to 2.10 hours) relative to the lightweight appearance codes. With multi-scale bilateral grid guided training, we found the rendering speed may improve due to the decreased number of unnecessary Gaussians.
@@ 5 Conclusion
We propose a multi-scale bilateral grid seamlessly integrating global appearance codes and pixel-wise bilateral grids to address photometric inconsistencies in autonomous driving neural rendering. Limitations. While our proposed multi-scale bilateral grid demonstrates significant improvements in handling photometric inconsistencies and enhancing geometric accuracy in driving scene reconstruction, certain limitations warrant discussion. (1) The computational remains higher than simpler appearance code techniques; this could be a factor in highly resource-constrained deployment scenarios. (2) while effective for many dynamic elements, modeling extremely complex, fast-moving, or highly non-rigid objects with complete fidelity remains an ongoing challenge.

### v4

@@ 4.1 Experimental Setup
Datasets.  We evaluate on four autonomous driving datasets: Waymo Sun et al. (2020), NuScenes Caesar et al. (2020), Argoverse Wilson et al. (2023), and PandaSet Xiao et al. (2021), selected for their diversity in sensor configurations (LiDAR, camera specifications), environmental conditions (lighting, seasons), and geographic locations. Identical model hyperparameters are used across all datasets to test generalization capability. Technical specifications including camera counts and ego-vehicle view cropping details are provided in Appendix A2.
Baseline.  Our focus is on modeling appearance variance across viewpoints, we build upon the open-source works DriveStudio Chen et al. (2024b), ChatSim Wei et al. (2024), StreetGS Yan et al. (2024a) and test three approaches: appearance codes, standalone bilateral grids, and our multi-scale bilateral grids. This allows for direct comparison of how each method handles photometric inconsistencies across viewpoints.
Table 2: Comparison with different baseline approaches for scene reconstruction. Performance is reported as the average over the NuScenes Caesar et al. (2020) scenarios.
@@
Implementation details.  We evaluate our method on many autonomous driving datasets to demonstrate its robustness. For each dataset, We train all the methods on a single NVIDIA L40 GPU and compare the 3D reconstruction and novel view synthesis (NVS) results between camera and LiDAR data. Additionally, we ablate key components of our method and quantify their impact on both appearance and geometry quality.
@@
Furthermore, we introduce two regularization terms to enhance image fidelity: Adaptive Total Variation Regularization. This term encourages smoothness and reduces noise while preserving image details, and Circle Regularization Loss. This loss applies inverse transformation to the ground-truth images, preventing discrepancies and image quality degradation. As visually demonstrated in Fig. 5, these terms effectively improve image fidelity. Detailed are provided in Appendix A1.
2) Rendering. For dynamic rendering using our multi-scale bilateral grid framework, we employ a specific interpolation strategy when encountering novel test images, especially those simulating dynamic ISP conditions. We first conducts a temporal proximity search to identify the most relevant grids. Following this, we perform scale-specific interpolation using the two nearest grids found.
Table 3: Evaluation on Challenging Scene Subsets. Results on systematically selected extreme cases. Our method shows substantially improvements when photometric inconsistencies are severe.
Scene Type Method Reconstruction Novel View Synthesis Geometry
Challenging Scenes (18 scenes) OmniRe $26.84$ $0.847$ 0.239 $24.46$ $0.758$ 0.263 $0.650$ $3.176$ $0.063$
Ours 28.13 0.856 0.223 25.17 0.763 0.250 0.523 3.120 0.023
Night Scenes (4 scenes) OmniRe $28.34$ $0.805$ 0.338 $26.54$ $0.753$ 0.360 $0.586$ $2.767$ $0.041$
Ours 29.13 0.812 0.330 27.21 0.762 0.349 0.500 2.731 0.016
Sun Flare Scenes (5 scenes) OmniRe $26.68$ $0.866$ 0.195 $24.98$ $0.796$ 0.212 $0.781$ $3.073$ $0.079$
Ours 28.65 0.879 0.173 25.84 0.803 0.192 0.593 2.994 0.022
Figure 3: Qualitative comparison across datasets. Our method versus baselines on Waymo, NuScenes, Argoverse, and PandaSet.
Figure 4: Photometric consistency improves geometry. Our framework (c) outperforms appearance codes (a) and single bilateral grids (b) by addressing optimization challenges and enhancing geometric modeling. This yields lower Chamfer Distance and fewer floaters in dynamic, decoupled driving scenes. Yellow indicates high LiDAR error, while Purple indicates low LiDAR error.
@@ 4.2 Quantitative Evaluation
Reconstruction and Geometric Quality. We present a comprehensive quantitative analysis in Tab. 1, comparing our full model against the OmniRe and its variants with appearance codes (w/AC) and a single bilateral grid (w/BG). Our method consistently sets a new state-of-the-art across all four datasets in both appearance and, most critically, geometric metrics.
In terms of appearance fidelity, our model achieves the highest PSNR and SSIM scores. For example, on NuScenes, our method reaches a PSNR of 27.69, a significant jump from 26.37 (OmniRe) and 26.38 (OmniRe w/AC). This demonstrates the superior ability of the multi-scale grid to model complex photometric variations.
However, the most striking result is the substantial improvement in geometric accuracy. Our method drastically reduces the Chamfer Distance, a key indicator of geometric fidelity. On the Waymo dataset, our model hits a CD of 0.989, outperforming the best baseline (OmniRe w/AC, 1.378) by a remarkable 28.2%. This trend holds across all datasets, with significant reductions in Chamfer Distance (CD), Root Mean Squared Error (RMSE), and Median Squared Depth Error (Depth). This result provides strong evidence for our central hypothesis: by accurately resolving photometric ambiguities, our multi-scale approach effectively eliminates geometric artifacts like "floaters" and produces a much cleaner, more accurate 3D reconstruction. The single bilateral grid (w/BG), despite its high expressiveness, often fails to converge and even degrades performance, highlighting the optimization challenges that our multi-scale, coarse-to-fine framework successfully overcomes.
NVS. As shown in Tab. 1, our model’s superior reconstruction quality translates directly to improved NVS. Our method consistently outperforms all baselines in NVS metrics across the four datasets. For instance, on PandaSet, we achieve a PSNR of 27.89, surpassing the next best (OmniRe w/AC at 27.51). This indicates that the cleaner geometry and more consistent appearance learned by our model generalize better to unseen viewpoints, producing higher-fidelity renderings.
Generalization to Other Methods. To demonstrate the broad applicability of our approach, we integrated our multi-scale bilateral grid into two other state-of-the-art methods, ChatSim Wei et al. (2024) and StreetGS Yan et al. (2024a). The results on NuScenes, presented in Tab. 2, are compelling. Our module brings substantial gains to both frameworks. For StreetGS, incorporating our method boosts the reconstruction PSNR from 25.74 to 27.90 and slashes the Chamfer Distance from 1.604 to 1.272. These significant improvements underscore that our multi-scale grid is not just a bespoke enhancement for one architecture but a generalizable and powerful module for improving photometric and geometric consistency in Gaussian Splatting-based scene reconstruction.
Performance on Challenging Scenarios. To validate robustness, we curated a subset of 18 challenging scenarios with extreme photometric inconsistencies. As shown in Table 3, our method’s advantages are substantially more pronounced here than on the standard datasets. Across this entire subset, our model achieves a +1.29 dB PSNR gain. This improvement is even more significant in specific difficult sub-categories; for instance, in "Sun Flare Scenes," the gain reaches +1.97 dB. This demonstrates our model’s strong effectiveness in resolving severe, real-world edge cases.
@@ 4.3 Qualitative Evaluation
Rendering Results.  Fig. 3 and Fig. 4 present visualization results of our method on scenes from Waymo Sun et al. (2020), NuScenes Caesar et al. (2020), Argoverse Wilson et al. (2023), and PandaSet Xiao et al. (2021), demonstrating its capability to handle diverse conditions that might otherwise lead to inconsistencies.
Figure 5: Ablation visualizations. Effects of circle regularization and adaptive total variation losses.
Figure 6: Affine transformation visualization. The first three rows show transformations and distributions for the first level, first two levels, and all three levels of our multi-scale approach. The final row compares against a single bilateral grid, illustrating a coarse-to-fine enhancement with richer transformation diversity.
@@
Table 4: Computational Efficiency Analysis. † denotes rendering with bilateral grid processing active.
Analysis of Learned Affine Transformations. To gain deeper insight into how our model functions, we analyze the distribution of the learned affine transformations in Fig. 6. The histograms for the original single-scale bilateral grid (BG) exhibit a distinct bi-modal, peaky distribution for each viewpoint. This suggests that the single grid learns a limited set of dominant photometric corrections for each view, struggling to capture the full spectrum of variations. In stark contrast, the aggregated histogram from our multi-scale grid (across all levels and viewpoints) is significantly flatter and more dispersed. This is a direct visualization of our framework’s core strength. The coarse-level grid first establishes a view-dependent baseline appearance, capturing the dominant global photometric shift (interestingly, its own histogram is also peaky, but varies across views). Then, the medium and fine grids learn residual transformations in a coarse-to-fine manner, correcting the errors from the preceding levels. This hierarchical, residual refinement process allows the model to represent a much broader and more diverse range of photometric transformations. The flatter histogram is empirical evidence of this enhanced representational power, which is crucial for achieving consistent, high-quality renderings across the diverse and challenging views encountered in autonomous driving.
@@ 4.4 Ablation Studies
Table 5: Ablations. Numbers in Guidance Map Resolution indicate downsample factors from coarse to fine level, and numbers in Grid Size Combinations indicate grid size.
Guidance Map Resolution. We test different downsampling factors for the guidance map at each grid level. The results in Tab. 5(a) show that a moderate downsampling of (2,2,1) from coarse to fine yields the best balance, achieving the highest NVS PSNR (24.64). Using the full resolution (1,1,1) slightly degrades performance, likely due to overfitting to noise, validating our patch-based aggregation strategy.
Grid Size Combinations. The combination of grid sizes is crucial. As shown in Tab. 5(b), removing the coarsest grid level (e.g., using only (4,4,2)+(8,8,4)) leads to a dramatic increase in CD, highlighting the importance of the coarse grid for establishing global geometric consistency. Crucially, a baseline with only a single fine-resolution grid (8,8,4) performs poorly, confirming that finer grids fail to converge without the stable, coarse-level initialization our method provides.
Loss Functions. Ablating our proposed regularization losses in Tab. 5(c) shows their contribution. Removing the circle regularization loss slightly degrades both reconstruction and NVS performance. The effect of the adaptive total variance loss is more subtle in metrics but visibly improves rendering quality by reducing noise, as shown in Fig. 5.
Efficiency. As detailed in Tab. 4, our method strikes an effective balance between performance and computational cost. The single large bilateral grid (w/BG) leads to a massive parameter count (27.8M) and the longest training time (2.85h). In contrast, our multi-scale approach requires only 3.9M parameters and a modest 9% increase in training time (2.10h vs. 1.93h) compared to the appearance code baseline, while delivering substantially better results. During inference, our method maintains high frame rates, demonstrating its practicality.
@@ 5 Conclusion
We introduced a novel multi-scale bilateral grid that unifies global appearance codes and pixel-wise bilateral grids into a single, hierarchical framework. Our extensive experiments demonstrate that this approach not only effectively models and corrects complex photometric inconsistencies in autonomous driving scenes but, more importantly, leads to a significant and crucial improvement in the geometric accuracy of the final reconstruction. By establishing this strong link between advanced appearance modeling and geometric fidelity, our work paves the way for more robust and reliable neural rendering systems in safety-critical, real-world applications.
Limitations. While our method shows significant improvements, some limitations remain. (1) The computational overhead, though much lower than a single large bilateral grid, is still higher than simple appearance codes, which could be a consideration for resource-constrained scenarios. (2) Modeling extremely fast-moving or highly non-rigid objects where LiDAR and camera data misalign remains an open challenge.
Appendix A Additional Implementation Details
A.1 Loss Functions and Optimization
We jointly optimize our multi-scale Gaussian scene representation by minimizing a composite loss function $\mathcal{L}_{\text{total}}$:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{recon}}+\lambda_{\text{TV}}\mathcal{L}_{\text{TV}}+\lambda_{\text{circle}}\mathcal{L}_{\text{circle}}$ (10)
where $\mathcal{L}_{\text{recon}}$ is the primary reconstruction loss, and $\mathcal{L}_{\text{TV}}$ and $\mathcal{L}_{\text{circle}}$ are regularization terms.
A.1.1 Reconstruction Loss
The core reconstruction loss, $\mathcal{L}_{\text{recon}}$, drives the accurate reproduction of both RGB and depth information:
$\mathcal{L}_{\text{recon}}=\lambda_{r}\mathcal{L}_{1}+(1-\lambda_{r})\mathcal{L}_{\text{SSIM}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{o}\mathcal{L}_{o}$ (11)
Here, $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$ measure the image-space difference. $\mathcal{L}_{d}$ represents the loss between the rendered depth and the ground truth LiDAR depth. $\mathcal{L}_{o}$ is an opacity regularization term that encourages alignment with a non-sky mask.
A.1.2 Adaptive Total Variation (TV) Regularization
To encourage smoothness and reduce noise in the optimized grid representations, we incorporate a TV regularization term, $\mathcal{L}_{\text{TV}}$, at each level $l$ of the multi-scale representation:
$\mathcal{L}_{\text{TV}}=\sum_{l}k^{(l)}\cdot\frac{1}{|\mathcal{A}^{(l)}|}\sum_{i,j,k}\sum_{\mathbb{D}\in\{x,y,z\}}\left\|\Delta_{\mathbb{D}}\mathcal{A}^{(l)}(i,j,k)\right\|_{2}^{2}$ (12)
The adaptive weight $k^{(l)}$ for each level $l$ is proportional to the grid size, applying stronger smoothing to finer, higher-resolution grids and lighter regularization to coarser grids.
$k^{(l)}=a\sqrt{H^{(l)}\cdot W^{(l)}\cdot D^{(l)}}+b$ (13)
A.1.3 Circle Regularization for Photometric Consistency
To constrain the noise introduced by photometric corrections and prevent overly aggressive alterations, we introduce a circle regularization loss, $\mathcal{L}_{\text{circle}}$:
$\mathcal{L}_{\text{circle}}=\sum_{(u,v)\in\mathcal{S}}\left\|I^{r}(u,v)-\bar{A}^{-1}(I^{gt}(u,v))\right\|_{2}^{2}$ (14)
where $\mathcal{S}$ denotes the set of pixel coordinates. This loss encourages the rendered image $I^{r}$ to be reconstructible from the ground truth image $I^{gt}$ via an inverse appearance transformation $\bar{A}^{-1}$.
A.1.4 Coarse-to-Fine Optimization Strategy
We employ a coarse-to-fine optimization strategy by utilizing level-dependent learning rates (e.g., $1\times 10^{-5},3\times 10^{-5},1\times 10^{-4}$ from coarse to fine). Coarser grids, assigned higher learning rates, rapidly learn the global scene illumination, while finer grids, with lower learning rates, hierarchically refine high-frequency photometric details. This staged optimization enhances stability.
A.2 Dynamic Rendering for Real-World ISP Adaptation
To bridge the domain gap between the fixed ISP used during training and dynamic real-world ISP pipelines encountered during inference, we employ an interpolation strategy. For a novel test image with timestamp $t_{\text{novel}}$, we proceed as follows:
  1. 1.
    Temporal Proximity Search: We identify the two training timestamps from the same camera, $t_{1}$ and $t_{2}$, that are temporally closest to $t_{\text{novel}}$ (such that $t_{1}\leq t_{\text{novel}}\leq t_{2}$).
  2. 2.
    Scale-Specific Grid Interpolation: We perform linear interpolation specifically on the coarse and medium-scale bilateral grids ($\mathcal{A}^{(0)}$, $\mathcal{A}^{(1)}$) derived from these two timestamps.
The interpolated grid $\hat{\mathcal{A}}^{(l)}$ for $l\in\{0,1\}$ is formulated as:
$\hat{\mathcal{A}}^{(l)}=\omega\mathcal{A}^{(l)}_{t_{1}}+(1-\omega)\mathcal{A}^{(l)}_{t_{2}}$ (15)
The temporal interpolation weight, $\omega$, is determined by proximity:
$\omega=\frac{t_{2}-t_{novel}}{t_{2}-t_{1}}$ (16)
The fine-scale grid ($l=2$) is not used during interpolation, as it is primarily intended to capture scene-intrinsic details rather than global ISP variations.
Appendix B Dataset Details
This section provides dataset-specific details regarding our evaluation protocol, including sequence IDs for reproducibility.
B.1 Waymo Open Dataset Sun et al. (2020)
We use all five cameras and all LiDAR sensors. We conduct our experiments on the following 6 sequences, selected according to Tonderski et al. (2024); Wei et al. (2024):
  • segment-10017090168044687777
  • segment-10061305430875486848
  • segment-10584247114982259878
  • segment-15090871771939393635
  • segment-4458730539804900192
  • segment-5835049423600303130
B.2 NuScenes Caesar et al. (2020)
We utilize all six available cameras and all LiDAR sensors. We select the following 8 sequences: 152, 164, 171, 200, 209, 359, 529, 916, which is an extension of the dataset used by Tonderski et al. (2024). To address ego-vehicle visibility, we crop the bottom 80 pixels from the back camera images.
B.3 PandaSet Xiao et al. (2021)
We utilize six cameras and one LiDAR unit. We specifically evaluate on the following 10 challenging nighttime sequences: 063, 066, 070, 073, 074, 077, 078, 079, 088, 149. To mitigate ego-vehicle artifacts, we apply a bottom crop of 260 pixels to the back camera images.
B.4 Argoverse2 Wilson et al. (2023)
We leverage the seven ring cameras and both LiDAR sensors available in the dataset. We conduct our experiments on the following 9 sequences, in line with Tonderski et al. (2024):
  • 05fa5048-f355-3274-b565-c0ddc547b315
  • 0b86f508-5df9-4a46-bc59-5b9536dbde9f
  • 185d3943-dd15-397a-8b2e-69cd86628fb7
  • 25e5c600-36fe-3245-9cc0-40ef91620c22
  • 27be7d34-ecb4-377b-8477-ccfd7cf4d0bc
  • 280269f9-6111-311d-b351-ce9f63f88c81
  • 2f2321d2-7912-3567-a789-25e46a145bda
  • 44adf4c4-6064-362f-94d3-323ed42cfda9
  • 5589de60-1727-3e3f-9423-33437fc5da4b
To minimize ego-vehicle interference, we apply a bottom crop of 250 pixels to the front center, rear left, and rear right camera views.
Appendix C Additional Quantitative Results
This section presents a comprehensive breakdown of our model’s performance, including per-scene metrics, integration results with SOTA methods, and robustness checks on challenging subsets.
C.1 Per-Scene Geometry and Appearance Evaluation
We provide detailed metrics for geometry (Tab. 6 - 9) and appearance (Tab. 10 - 13).
C.2 Integration Evaluation with SOTA Methods
Tab. 14 and 15 provide comparisons with SOTA methods like ChatSim Wei et al. (2024) and StreetGS Yan et al. (2024a) on the NuScenes dataset.
C.3 Performance on Challenging Scenarios
To validate robustness, we curated subsets of challenging scenarios with extreme photometric inconsistencies (e.g., night scenes, sun glare, rain, and complex reflections). As shown in Tables 16, 17, and 18, our method’s performance gap widens significantly compared to all baselines (OmniRe, w/ AC, w/ BG) in these difficult conditions, demonstrating the superior robustness of our multi-scale design.
  • Challenging Argoverse (6 scenes): 49e970c4 (Extreme Lighting, Overexposure, Lens Flare), 8184872e (Complex Shadows, Direct Sunlight), 91923e20 (Extreme Lighting, Camera Artifacts), a8a2fbc2 (Low-Angle Sun), b403f8a3 (Overpass Structure, Extreme Lighting Transitions), eaaf5ad3 (Multiple Artificial Lights, Wet Surfaces, Rainy Conditions).
  • Challenging Waymo (6 scenes): segment-10 (Night Scene, Artificial Lighting, Lens Flare), segment-30 (Night Scene, Rainy Conditions, Specular Reflections), segment-539 (High Dynamic Range, Low-Angle Sun, Sun Glare), segment-550 (High Dynamic Range, Storefront Reflections, Hard Shadows), segment-561 (Low-Angle Sun, Lens Flare, Overexposure), segment-570 (Rainy Conditions, Low Light, Motion Blur).
  • Challenging Pandaset (6 scenes): 19, 21, 29, 48, 52 (all Low-Angle Sun with Hard/Deep Shadows or High Dynamic Range), 63 (Night Scene, Multiple Light Sources, Glare).
C.4 Vehicle-Specific Metrics
To specifically evaluate performance on challenging, highly reflective surfaces such as car windows, we used semantic masks to isolate vehicles. Table 19 shows that our method improves not only appearance (PSNR) but also, crucially, geometry (CD) for these specific objects. This supports our claim that our method mitigates the negative geometric impact of such severe view-dependent effects.